# Learning From Software Failures: A Case Study at a National Space Research Center

### v2

@@ 6.4. Future directions
Based on the current practices (cf. section 5.1), their flaws (cf. section 5.2), and their challenges (cf. section 5.3) in learning from failures that we observed in our case study, we propose three opportunities for future research.
First, more research is needed to evaluate the impact of different failure learning practices, particularly the tradeoffs between lightweight, more ad hoc methods and structured approaches. When is informal reflection enough? When does structure pay off? What contextual factors (e.g., team size, mission type) make one approach preferable over another?
Second, our data suggest the need to design and evaluate lightweight failure learning processes that support gathering, documentation, and reflection of lessons without imposing significant overhead. As participants repeatedly noted, time and resources are major constraints, even when the cost of failures (i.e., the value of failure learning) is acknowledged. Milestone-based postmortems or team-driven reflection sessions may be more sustainable than rigid weekly reviews or top-down documentation mandates.
Third, automation tools may offer a promising avenue to reduce the burden of capturing and retrieving failure knowledge. There are existing tools to help manage incidents (Fir, [n. d.]; PagerDuty, 2023), but they are limited help for leveraging lessons learned. Recent works have explored using LLMs to synthesize incident data, from various sources such as public reports (Anandayuvaraj et al., 2024) and logs (Ahmed et al., 2023) into postmortem reports, and to apply them to understand (Jin et al., 2023), conduct root cause analysis (Zhang et al., 2024), and suggest mitigation (Ahmed et al., 2023) for incidents. Combining such automated reports, with Retrieval-Augmented Generation (RAG) (Lewis et al., 2020) techniques may also enable engineers to easily leverage failure data and lessons. For example, this could enable a failure-aware chatbot to provide engineers with targeted, context-sensitive advice grounded in past failures. However, as our findings highlight, existing tools such as wikis and questionnaires are often underutilized, suggesting that automation efforts must integrate seamlessly into developers’ existing workflows to encourage adoption.
@@ 7. Conclusion
Learning from failures is widely regarded as essential for improving reliability in high-stakes domains, however there is a knowledge gap on how it is done in practice. This study presents the first in-depth qualitative analysis of how software engineers in an HRO gather, document, share, and apply lessons from failures. Through interviews with 10 engineers at a space research center and 5 pilot interviews at other HROs, we found that failure learning is informal, ad-hoc, and inconsistently integrated into software development practices, even in settings where reliability is critical. Our findings reveal challenges including time and resource constraints, knowledge loss due to team turnover and fragmented documentation, and weak process enforcement. Although HROs and IT organizations are held as exemplars for structured postmortems, the feedback loop from documented insights to software process improvements remains poorly understood across domains.
Our findings shed light on the real-world practices, challenges, and organizational factors shaping failure learning practices within software teams in HROs, and suggests directions for designing better processes, tooling, and organizational structures to support failure-aware software development life-cycle.

### v3

@@ 6.4. Future Directions
Based on the current practices ( section 5.1), their flaws ( section 5.2), and their challenges ( section 5.3) in learning from failures observed in our case study, we propose two opportunities for future research.
First, more research is needed to evaluate and measure the value added by different failure learning practices, particularly the tradeoffs between lightweight ad-hoc methods and structured approaches. When is informal reflection enough? When does structure pay off? What contextual factors (e.g., team size, mission type) make one approach preferable over another? We note that there are no established ways to measure the value or effectiveness of such practices, making this measurement gap itself an open research problem. Additionally, all subjects repeatedly noted that time and resources are major constraints (Kasi et al., 2008; Desouza et al., 2005; Vieira et al., 2019; Lyytinen and Robey, 1999; Dingsøyr, 2005), even when the cost of failures (i.e., the value of failure learning) is acknowledged. Our data suggest the need to design and evaluate lightweight failure learning practices that support gathering, documentation, and reflection of lessons without imposing significant overhead. Although this need has been discussed in prior works (Dingsøyr et al., 2001; Dingsøyr, 2005; Myllyaho et al., 2004), it appears to still be unresolved in practice.
Second, automation tools may reduce the burden of capturing and retrieving failure knowledge. There are existing tools to help manage incidents (Fir, [n. d.]; PagerDuty, 2023), but they are limited help for leveraging lessons learned. Recent works have explored using LLMs to synthesize incident data into reports (Anandayuvaraj et al., 2024; Jin et al., 2023). Combining such automated reports with Retrieval-Augmented Generation (RAG) (Lewis et al., 2020), or other Generative Artificial Intelligence (GenAI) (Alavi et al., 2024) techniques, may also enable engineers to easily leverage organizational knowledge on failure data (Ahmed et al., 2023) and lessons. For example, this could enable a failure-aware chatbot to provide engineers with targeted, context-sensitive advice grounded in past failures. However, as our findings highlight, existing tools such as wikis and questionnaires are often underutilized, suggesting that automation efforts must integrate seamlessly into developers’ existing workflows to encourage adoption.
@@ 7. Conclusion
Learning from failures is widely regarded as essential for improving reliability in high-reliability organizations, however there is a knowledge gap on how it is done in practice. This study presents the first in-depth qualitative analysis of how software engineers in an HRO gather, document, share, and apply lessons from failures. Through interviews with 10 engineers at a space research center, supplemented by 5 pilot interviews at other HROs, we found that failure learning is informal, ad-hoc, and inconsistently integrated into software development practices, even in settings where reliability is critical. Our findings reveal challenges including time and resource constraints, knowledge loss due to team turnover and fragmented documentation, and weak process enforcement. Additionally, the feedback loop from lessons learned to improvements remains poorly understood . Our findings shed light on the real-world practices and challenges for learning for software failures in HROs, and suggests future research directions to support a failure-aware software development life-cycle.
8. Acknowledgments
This work was supported by the German DAAD Rise fellowship, the German Aerospace Center (DLR), Cummins, the Qualcomm Innovation Fellowship, and the Jet Propulsion Laboratory  [^3] Portions of this work were performed by the Jet Propulsion Laboratory, California Institute of Technology, under a contract with the National Aeronautics and Space Administration (80NM0018D0004). Reference herein to any product, process, or service by trade name, trademark, or manufacturer, does not imply its endorsement by the United States Government or JPL.
9. Data Availability
Our artifact (https://doi.org/10.5281/zenodo.17754815) contains our interview protocol, codes, and themes. We omit transcripts due to the difficulty in anonymizing a case study and the sensitive nature of the data.